OpenStudy (anonymous):

SAT question: If (a+b)^2 = 64 and a^2+b^2 = 40, what is the value of 4ab?

OpenStudy (anonymous):

\[(a+b)^2=a^2+b^2+2ab=64\] and \[a^2+b^2=40\] can work it out now right?